For 2019 Annita has gross income of $100,000 and deduction of $160,000 for regular tax purposes.

  1. What are the tax consequences to Annita in the current year?

  2. Would the result have any impact on Annita’s tax liability in the prior year if it had taxable income in that year of $50,000?

  3. Suppose that for the following year Annita has taxable income from all activities in that year of 30,000. Assume the tax rate is a flat 21% of taxable income. How much federal tax will Bagco pay in that year?
Answer #1

Solution:-

According to tax laws, In 2019 Annita can carry forward 60000$ (160000-100000) to following years .Annita dont have to pay any tax in the current year.She can adjust her carry forward losses with subsequents years income.

According to tax law,tax credit or losses cant be adjusted with prior periods tax liablity.Prior years tax liablity will be already paid in the previous year so that it can be refunded.

From Dec 2017,One can adjust only 80% of the total income with the carry forward losses.So Annita has to pay tax for the remaining amount i.e 30000*20%=6000$

Tax liablity=6000*21%=1260$